Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
pinto, ruggi - 0:50 Friday 01 March 2024 (63452) Print this report
Vertical and Roll MIR resonance damper + PR TX control filter test

Today's activity was delayed by the presence of high wind activity up untill around 16.00 UTC.

Over the shift we experienced several unlocks: a couple at CARM null 1f due to a low gain of DIFFp TY loop, which was affected by a low frequency oscillation (0.4 Hz) during the transition from CARM null 3f to 1f (Diffp drift control -> full bandwidth). Almost doubling the loop gain cured the issue.

A couple of other unlocks occurred in the steady state at LowNoise 3 after few minutes of lock, right after the glitch, with the arise of a 1.1 Hz oscillation, which we think it may be related to the PR TX loop.


Regarding the activity of the shift we performed several test by restoring former damper loops of the MIR roll and vertical modes of the four test masses: 9.8 Hz (TZ roll) and 7 Hz (Y vertical).

Better details will be posted in an other entry. Here we reported the GPS time stamps of the tests:

initial test with roll damper:
- 16.37.20 UTC TZ damper of NI switched on in LN2.

We encountered some issue with the reload of some old version of the configuration file ITF_LOCK.ini.  After we restored the correct locking parameters we were able to reach LowNoise 3 but we unlocked shortly after.

test with vertical damper:
- 21.23 UTC:  reference data with BNS range stable higher than 53 Mpc in LN3.
- 21.26.30 UTC: we reduced the dither lines on WI, NI, WE, NE by a factor 5, from 0.005 to 0.001. 
- 21.32.20 UTC: switched on the NI WI NE WE 7 Hz vertical damper. The effect on DARM spectra was immediatly visible as the 7 Hz resonance decreased in amplitude.

test with roll damper:
- 21.42.30 UTC: we switched only the loops on WI and NI, since in a previous test the WE test caused an unlock.
- 21.55.45 UTC we switched on the NE damper loop, but shortly after we opened all three loops, since the mode was somehow excited.

In order to understand if the NE damper is somehow affected or can lead to some instability we tried to engage the loop alone with all the other 3 open.
- 22.07 UTC: test with NE standalone, which worked quite fine and it didn't seem to be unstable.
- 22-11.50 UTC: same test with the WE loop: we didn't unlock but the loop was not behaving well since an increase of lf correction was noticed.


PR TX control test:
as a last test,  at 22.13.45 UTC we engaged the PR TX control filter, which was tested once some time ago. This filter is aimed to improve of almost a factor 2 the rms of the loop, and to reduce the amplitude of the sbs on several peaks on Hrec due to non linear couplings. The filter behaved well and it doesn't seem to suffer low gain instability at around 1.1 Hz.
In the future it seems worthed to evaluate the stability of the loop at every step of the acquisition in order to leave this one as default.
At 22.37 UTC we reverted the control filter to the standard one.

We left the ITF locked in LN3, with autorelock engaged and with all the damper loops OPEN, but with the amplitude of the dither lines reduced of a factor 5.

Comments to this report:
pinto, ruggi - 11:39 Friday 01 March 2024 (63454) Print this report

Even if we tought we reverted the nominal control filter for PR TX, this morning we found that the new one was still running.

At a first sight, by checking the data, it looks like it worked fine for the whole acquisition. Maybe it should be good to check the right gain values for the different steps.

We left the new TX controller as default.

Search Help